# A Synthetic Benchmark for Collaborative 3D Semantic Occupancy Prediction in V2X Autonomous Driving

### v1

@@ Introduction
Table 2: 3D semantic occupancy prediction results on our Co3SOP dataset. We report the mIoU of some state-of-art methods and our baseline model under three tasks with different ranges, i.e., $25.6\times 25.6\times 4.8m^{3}$, $51.2\times 51.2\times 4.8m^{3}$ and $76.8\times 76.8\times 4.8m^{3}$, as well as the IoUs for some object class. The top performances are highlighted in bold.
@@
Image Backbone. We adopt ResNet101-DCN (He et al. 2016) as the image backbone to extract multi-scale image features from multi-view camera images. The extracted features are further refined using a Feature Pyramid Network (FPN) (Lin et al. 2017) to aggregate contextual information across scales.
Image Deformable Cross-Attention. To integrate the 2D image feature into the 3D space, we apply a image deformable cross-attention as in SurroundOcc (Wei et al. 2023). This image deformable cross-attention mechanism aggregates multi-view 2D features and learns to map them into the corresponding 3D voxel locations, by considering geometric relationships, ensuring that spatial consistency is maintained when transitioning from 2D to 3D space.
3D Affine Transformation. Before multi-agent fusion, voxel features from neighboring agents are transformed into the ego vehicle’s coordinate frame using a 3D affine transformation module inspired by Spatial Transformer Networks (Jaderberg et al. 2015). The transformation matrix is computed from relative poses between agents, ensuring that voxel features from different coordinates are properly aligned for subsequent attention-based fusion.
Mask Voxel Deformable Cross-Attention. To enable robust and adaptive fusion, each vehicle maintains a learnable confidence mask that estimates the voxel-wise reliability of its features. In addition, a binary warping mask is generated during feature alignment to indicate valid projection areas. These two masks are combined to form a hybrid attention mask, which modulates the voxel deformable cross-attention. This design allows the ego vehicle to selectively aggregate features from its collaborators based on both alignment accuracy and confidence, suppressing unreliable or noisy regions. The voxel deformable attention dynamically samples neighboring points across agents, guided by learned offsets and the hybrid attention mask.
Loss
To supervise collaborative 3D occupancy prediction, we combine a voxel-wise cross-entropy loss for semantic classification, a scene-class affinity loss (Cao and de Charette 2022) to encourage consistency between semantic and geometric features, and a confidence loss that regularizes the learned confidence mask for inter-agent feature consistency.
Experiments
Dataset and Metrics
To evaluate the performance of collaborative 3D semantic occupancy prediction under varying spatial conditions, we define three perception range settings in our benchmark: $25.6\times 25.6\times 4.8\,m^{3}$ range with $0.1\,m$ voxel size, $51.2\times 51.2\times 4.8\,m^{3}$ range with $0.2\,m$ voxel size, and $76.8\times 76.8\times 4.8\,m^{3}$ range with $0.3\,m$ voxel size. These settings allow us to assess how collaborative perception improves scene understanding as the coverage area increases, particularly in occluded or distant regions where a single vehicle’s sensing capability is limited. For evaluation, we adopt Intersection-over-Union (mIoU), including IoU for each class individually, as well as the mean IoU (mIoU) across all semantic categories.
Benchmark Methods
@@
To establish a fair comparison under our benchmark, we evaluate four representative 3D semantic occupancy prediction models from two modalities. For LiDAR-based methods, we select SSCNet (Song et al. 2017) and LMSCNet (Roldão, de Charette, and Verroust-Blondet 2020); for camera-based methods, we include SurroundOcc (Wei et al. 2023) and OccFormer (Zhang, Zhu, and Du 2023). These models are chosen for their widespread adoption and strong performance in prior benchmarks. We adopt their official implementations and default configurations, modifying only the data loaders and batch sizes to ensure compatibility with the Co3SOP dataset.
Implementation Details
@@
For the image backbone, we adapt the weights from FCOS3D (Wang et al. 2021) as the pretrained weights and set the number of output levels to 4. And we then fed the image features of stage 1, 2, 3 into FPN and obtain 4 levels multi-scale image features. For the image and voxel deformable attention, we set the number of layers as 3 and 1. For intermediate voxel features, we apply a downsampling rate of $\frac{1}{4}$ and upsample them in the prediction head. For the collaborative setting, we set the maximum number of collaborating agents to 6 per ego vehicle, based on number of vehicles in OPV2V. During the training process, we apply multi-scale supervision for the multi-scale outputs from prediction head and image augmentation as in SurroundOcc (Wei et al. 2023). All experiments are conducted on 8 RTX 4090 GPUs.
@@
Table 3: Ablation study on the impact of collaboration in 3D semantic occupancy prediction.
Range Collaboration  Buldings  Fences  Other  Poles  Roadlines Roads Sidewalks  Vegetation  Vehicles  Walls  Trafficsigns  Ground  Bridge  Guardrail  Trafficlight  Terrain mIoU
@@
Benchmark Analysis
The performance of the benchmark methods, SSCNet, LMSCNet, SurroundOcc, and OccFormer, along with our proposed Co3SOP-Base, is presented in table 2, which reports mean IoU (mIoU) as well as class-wise IoUs across three perception ranges.
Among the four selected methods, OccFormer achieves the highest mIoU in the $25.6,m$ range (25.41), while SurroundOcc performs best in the $51.2,m$ and $76.8,m$ ranges (25.53 and 24.18, respectively). Notably, both camera-based methods consistently outperform LiDAR-based methods in overall mIoU, highlighting their advantage in capturing rich semantic features from dense RGB inputs. However, LiDAR-based models still demonstrate strengths in certain object classes particularly in short-range setting such as Vehicles and Poles. These categories benefit from LiDAR’s precise geometric structure. But LiDAR-based methods also struggle with detecting low-elevation or textureless categories such as roadlines, where geometric features are weak or absent.
Compared with the four selected methods, our proposed Co3SOP-Base consistently achieves the best overall performance across all three perception ranges, with mIoUs of 30.04 ($25.6m$), 27.50 ($51.2m$), and 27.00 ($76.8m$). These results highlight the effectiveness of collaborative voxel fusion, especially as the spatial extent increases.
Ablation Study
Effect of Collaborative Feature Fusion. We investigate the impact of inter-agent collaboration by comparing performance with and without V2V feature fusion across all three perception ranges. As shown in table 3, collaboration consistently improves mIoU across three ranges. The gains are most significant in the large-range setting, where agents have broader visibility overlap and benefit more from complementary viewpoints. Object-level analysis reveals that collaboration notably enhances performance for large or occlusion-prone classes such as Vehicles (e.g., $50.9834.43\xrightarrow{}$ 50.98 IoU at $76.8m$) and Poles (e.g., $16.1610.34\xrightarrow{}$ 16.16 IoU at $51.2m$). However, for thin or rare classes like Trafficlight, collaboration yields little to no improvement—and in some cases even slight drops (e.g., Trafficlight: $3.303.35\xrightarrow{}$ 3.30 at $51.2m$).
@@
Table 4: Ablation study on the impact of pose noise in collaborative 3D semantic occupancy prediction.
Robustness on Pose Noise We further assess the robustness of Co3SOP-Base under localization uncertainty by injecting Gaussian pose noise into the relative transformations used for inter-agent feature alignment. As shown in table 4, increasing the mean offset from $\mu=0.1m$ to $0.5m$ (with fixed standard deviation $\sigma=0.02$) leads to a gradual mIoU degradation across all ranges. This decline illustrates that misalignment in inter-agent transformations can affect collaborative feature quality. However, the performance remains relatively stable under moderate noise, suggesting that our warping-based spatial alignment and confidence-aware fusion retain reasonable robustness in the presence of pose perturbations.
Conclusion
We presented Co3SOP, a synthetic benchmark for collaborative 3D semantic occupancy prediction, built upon high-fidelity simulation and comprehensive voxel annotations. Unlike prior datasets limited by LiDAR sparsity, Co3SOP leverages a custom-designed voxel annotation pipeline to provide complete and fine-grained semantic voxel labels, enabling robust evaluation of multi-agent perception systems. We further proposed Co3SOP-Base, a baseline framework incorporating confidence and alignment aware masked voxel deformable attention for multi-agent feature fusion. Extensive experiments demonstrate the effectiveness of V2V collaboration in improving perception performance, especially in distant or occluded regions. The proposed Co3SOP serves as a dedicated resource for advancing research in voxel-level 3D scene understanding under collaborative scenarios.

### v2

@@ V-A Dataset and Metrics
TABLE II: 3D semantic occupancy prediction results on our Co3SOP dataset. We report the mIoU of some state-of-art methods and our baseline model under three tasks with different ranges, i.e., $25.6\times 25.6\times 4.8m^{3}$, $51.2\times 51.2\times 4.8m^{3}$ and $76.8\times 76.8\times 4.8m^{3}$, as well as the IoUs for some object class. The top performances are highlighted in bold.
@@
V-B Benchmark Methods
@@
To establish a fair comparison under our benchmark, we evaluate four representative 3D semantic occupancy prediction models from two modalities. For LiDAR-based methods, we select SSCNet [14] and LMSCNet [15]; for camera-based methods, we include SurroundOcc [35] and OccFormer [20]. These models are chosen for their widespread adoption and strong performance in prior benchmarks. We adopt their official implementations and default configurations, modifying only the data loaders and batch sizes to ensure compatibility with the Co3SOP dataset.
V-C Implementation Details
@@
For the image backbone, we adapted the weights from FCOS3D [39] as the pretrained weights and set the number of output levels to 4. And these image features from stages 1, 2, and 3 are then fed into the FPN to obtain 4 levels of multi-scale features. For the image and voxel deformable attention, we set the number of layers as 3 and 1. For intermediate voxel features, we apply a downsampling rate of $\frac{1}{4}$ and upsample them in the prediction head. For the collaborative setting, we set the maximum number of collaborating agents to 6 per ego vehicle, based on number of vehicles in OPV2V. During the training process, we apply multi-scale supervision for the multi-scale outputs from prediction head and image augmentation as in SurroundOcc [35]. All experiments are conducted on 8 RTX 4090 GPUs.
V-D Benchmark Analysis
The performance of the benchmark methods, SSCNet, LMSCNet, SurroundOcc, and OccFormer, along with our proposed Co3SOP-Base, is presented in table II, which reports mean IoU (mIoU) as well as class-wise IoUs across three perception ranges.
Among the four selected methods, OccFormer achieves the highest mIoU in the $25.6,m$ range (29.48), while SurroundOcc performs best in the $51.2,m$ and $76.8,m$ ranges (25.76 and 24.68, respectively). Notably, both camera-based methods consistently outperform LiDAR-based methods in overall mIoU. However, LiDAR-based models still demonstrate strengths in certain object classes particularly in small-range such as Vehicles and Poles.
Compared with the four selected methods, our proposed Co3SOP-Base consistently achieves the best overall performance across all three perception ranges, with mIoUs of 30.04 ($25.6m$), 27.50 ($51.2m$), and 27.00 ($76.8m$). Compared to the single vehicle methods, our proposed baseline gaines greater improvements as the prediction range increases, especially for object classes such as vehicles, guardrail and roadlines. These results highlight the effectiveness of collaborative voxel fusion, especially as the spatial extent increases.
@@ V-D Benchmark Analysis
TABLE III: Ablation study on the impact of collaboration in 3D semantic occupancy prediction.
Range
Collaboration
 Buildings
 Fences
 Other
 Poles
 Roadlines
Roads
Sidewalks
 Vegetation
 Vehicles
 Walls
 Trafficsigns
 Ground
 Bridge
 Guardrail
 Trafficlight
 Terrain
mIoU
@@
V-E Ablation Study
Effect of Collaborative Feature Fusion. We investigate the impact of inter-agent collaboration by comparing performance with and without V2V feature fusion across all three perception ranges. As shown in table III, collaboration consistently improves mIoU across three ranges. The gains are most significant in the large-range setting, where agents have broader visibility overlap and benefit more from complementary viewpoints. Object-level analysis reveals that collaboration notably enhances performance for large or occlusion-prone classes such as Vehicles (e.g., $34.43\xrightarrow{}50.98$ IoU at $76.8m$) and Poles (e.g., $10.34\xrightarrow{}16.16$ IoU at $51.2m$). However, for thin or rare classes like Trafficlight, collaboration yields little to no improvement—and in some cases even slight drops (e.g., Trafficlight: $3.35\xrightarrow{}3.30$ at $51.2m$).
@@ V-E Ablation Study
Robustness on Pose Noise We further assess the robustness of Co3SOP-Base under localization uncertainty by injecting Gaussian pose noise into the relative transformations used for inter-agent feature alignment. As shown in table IV, increasing the mean offset from $\mu=0.1m$ to $0.5m$ (with fixed standard deviation $\sigma=0.02$) leads to a gradual mIoU degradation across all ranges. This decline illustrates that misalignment in inter-agent transformations can affect collaborative feature quality. However, the performance remains relatively stable under moderate noise, suggesting that our warping-based spatial alignment and confidence-aware fusion retain reasonable robustness in the presence of pose perturbations.
TABLE IV: Ablation study on the impact of pose noise on collaborative 3D semantic occupancy prediction.
@@
VI Conclusion
We presented Co3SOP, a synthetic benchmark for collaborative 3D semantic occupancy prediction, built upon high-fidelity simulation and comprehensive voxel annotations. Unlike prior datasets which are limited by LiDAR sparsity, Co3SOP leverages a custom-designed voxel annotation pipeline to provide complete and fine-grained semantic voxel labels, enabling robust evaluation of multi-agent perception systems. We further proposed Co3SOP-Base, a baseline framework incorporating confidence and alignment aware masked voxel deformable attention for multi-agent feature fusion. Extensive experiments demonstrate the effectiveness of V2V collaboration in improving perception performance. This work addresses the critical gap in collaborative 3D voxel-level understanding by providing both a benchmark and a baseline, paving the way for safer and more robust autonomous driving systems.